OpenStudy (juan1857):

Let f(x)=-2x-7 and g(x)=-4x+3. Find (fog)(-5)

OpenStudy (littleelf):

You need to put the g function in f and then substitute the value of x with 3. ie. f(g(x)) = 2(-3x -7) -6 fog= -6x -20 fog(3)= -6*3 -20 fog(3)= -38 2. log(5x + 14) =0 take the log in RHS (5x + 14)= e^0 5x + 14= 1 => x= -13/5 Hope this helps.

OpenStudy (littleelf):

:P

OpenStudy (littleelf):

I hope I helped

OpenStudy (juan1857):

let me just process this, lol

OpenStudy (littleelf):

Lol

OpenStudy (juan1857):

wait, where did you get the 3?

OpenStudy (littleelf):

What 3

OpenStudy (juan1857):

the three here: You need to put the g function in f and then substitute the value of x with 3.

OpenStudy (littleelf):

I have this in my notes I have no Idea

OpenStudy (littleelf):

Its been a while

OpenStudy (juan1857):

oh lol

OpenStudy (littleelf):

Lol

OpenStudy (littleelf):

Start with the given property Plug in f%28x%29=x%5E2-6x%2B9 and g%28x%29=x-3 Group like terms Combine like terms So

OpenStudy (littleelf):

oops

OpenStudy (littleelf):

Here look http://prntscr.com/b9o36f

OpenStudy (juan1857):

lol

OpenStudy (littleelf):

That might help

OpenStudy (juan1857):

um this is a different type of question

OpenStudy (littleelf):

its another one to see

OpenStudy (littleelf):

to kinda help a little more

OpenStudy (juan1857):

it is (f-g)

OpenStudy (littleelf):

Opps I cut something outon my 1 answer

OpenStudy (littleelf):

o means (of) i.e. f of g i.e. f(g(x)). Here you go. You need to put the g function in f and then substitute the value of x with 3. ie. f(g(x)) = 2(-3x -7) -6 fog= -6x -20 fog(3)= -6*3 -20 fog(3)= -38 2. log(5x + 14) =0 take the log in RHS (5x + 14)= e^0 5x + 14= 1 => x= -13/5 Hope this helps.

OpenStudy (littleelf):

Ok

OpenStudy (juan1857):

i'm still kinda lost, but thanks anyways

OpenStudy (littleelf):

Im sorry I could not help u a lot

OpenStudy (juan1857):

its ok, at least you tried :)

OpenStudy (littleelf):

Ya I guess

OpenStudy (littleelf):

Make sure who ever helps u fully gets a medal!

OpenStudy (juan1857):

I wanna give you a medal for your kindness

OpenStudy (sshayer):

\[\left( fog \right)x=f \left\{ g \left( x \right) \right\}=f \left( -4x+3 \right)=-2\left( -4x+3 \right)-7=8x-13\] \[\left( fog \right)\left( -5 \right)=8*-5-13=-53\]

OpenStudy (littleelf):

aww ty @juan1857

OpenStudy (juan1857):

oh so do we practically just insert either g or f as x for the other?

OpenStudy (littleelf):

Yep

OpenStudy (juan1857):

ahh! Now I see

OpenStudy (littleelf):

Dont forget to medal whoever u think earned it

OpenStudy (juan1857):

I won't :)

OpenStudy (littleelf):

Ok :)

OpenStudy (littleelf):

ty

OpenStudy (juan1857):

my pleasure